# Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning

### v2

@@ 4.2. Greedy Prompt Search
Despite Gemini’s capabilities, its generated prompts may still be suboptimal in terms of triggering maximal toxicity. To this end, we propose a greedy prompt search algorithm to iteratively refine image prompts based on observed toxicity feedback. At each iteration, Gemini generates a new candidate prompt, which is used to synthesize an image via the diffusion model. This image, paired with the fixed text prefix, is input to the target VLM, and the resulting continuation is scored using a toxicity classifier (e.g., Detoxify (Hanu and Unitary team, 2020)). If the toxicity score increases relative to the previous iteration, the new image prompt is retained and the search continues; otherwise, the process terminates. This loop yields near-optimal image prompts that serve as strong initialization points for downstream adversarial image generation.
4.3. Reward Functions for RTD
A contribution of RTD lies in its dual-reward optimization, which explicitly balances attack effectiveness and semantic plausibility. We introduce two complementary rewards: a toxicity reward that drives harmful continuation, and an alignment reward that ensures prompt-consistent image generation.
Toxicity Reward: In our toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, paired with an image $\mathbf{x}$ generated by the RTD. The toxicity of the generated continuation is then assessed using the Detoxify classifier (Hanu and Unitary team, 2020), which evaluates six toxicity-related attributes (e.g., insult, threat, identity attack), each scored between 0 (least toxic) and 1 (most toxic). We use the average of these scores as the toxicity reward to guide diffusion-based adversarial image synthesis:
@@ 4.3. Reward Functions for RTD
$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$
@@
Alignment Reward: To encourage semantic alignment between the input prompt $p$ and the VLM’s interpretation of the generated image, we introduce an alignment reward that regularizes the image generation process. Following (Black et al., 2023), we use BERTScore (Zhang et al., 2019) to measure the similarity between the image prompt and the caption produced by the victim VLM. Specifically, the alignment reward is defined as:
@@
$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big{(}S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big{)}$
[…]
@@
$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$
This combined reward allows the RTD to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
Algorithm 1 Training Procedure of Red Team Diffuser
@@
23:  return Fine-tuned Diffuser $D_{\theta}^{*}$.
5. Experiments
@@
Datasets We use the RealToxicityPrompts benchmark (Gehman et al., 2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2024). We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
@@
Metric: (1) Toxicity Rate. To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we use the Detoxify classifier (Hanu and Unitary team, 2020) during training, and both Detoxify and the Perspective API during testing. These classifiers assess toxicity across six predefined attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). For evaluation, we compute the proportion of continuations whose toxicity score exceeds a threshold of 0.5 for each attribute. The use of the Perspective API at test time allows us to assess cross-classifier generalization and mitigate risks of reward hacking or classifier-specific bias. (2) CLIPScore. To evaluate the semantic alignment between generated images and their prompts, we use CLIPScore (Hessel et al., 2021), a reference-free metric based on the cosine similarity between CLIP-encoded image and text features. Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following (Hessel et al., 2021)) for alignment with human judgment. In our setting, CLIPScore serves as an auxiliary metric to verify that adversarial images preserve semantic plausibility without compromising visual coherence.
@@
Implementation Details We use Stable Diffusion v1.5 (Rombach et al., 2022) as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation. During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU. We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
5.1. Main Results
To evaluate the effectiveness of Red Team Diffuser (RTD), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM. We compare RTD against several baseline strategies to assess its advantage in inducing harmful outputs. We first include a Text-only Attack, in which the model is prompted solely with toxic text prefixes without any accompanying image. This serves as a unimodal baseline to quantify the additional effect of visual inputs on toxicity. Next, we define the Single-Prompt Attack, where Gemini generates a single image prompt per prefix without any optimization. This provides a naive multimodal baseline, capturing the effect of directly pairing a non-optimized image with a toxic prefix. The Greedy-Prompt Attack improves upon this by iteratively refining the image prompt using our proposed greedy search algorithm. Finally, we evaluate the full RTD framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards. We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
@@ 5.1. Main Results
Table 1. Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
@@
RTD w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
RTD w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
RTD w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
RTD w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
@@
RTD w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
RTD w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
RTD w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
RTD w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
Table 2. Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@
RTD w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
RTD w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
RTD w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
RTD w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
@@
RTD w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
RTD w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
RTD w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
RTD w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
@@
Original Attack Set Table 1 presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the generated outputs, with ‘Any*’ indicating whether a continuation exhibits at least one of the six toxic attributes. We begin by examining the Text-only Attack, where only the toxic text prefix is provided without any visual input. Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate—achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective. This indicates that current alignment mechanisms in VLMs remain insufficient for preventing harmful completions even in unimodal settings, exposing a fundamental vulnerability in their default safety alignment. When introducing visual input, the Single-Prompt Attack slightly alters the toxicity pattern but does not significantly improve attack effectiveness, serving as a weak multimodal baseline. In contrast, the Greedy-Prompt Attack significantly amplifies toxicity. For instance, using the Detoxify classifier, the ‘Any*’ rate increases from 55.72% (Single-Prompt) to 64.12%, illustrating the benefit of greedy prompt search in finding high-risk image prompts. The full Red Team Diffuser (RTD) further enhances attack strength. RTD w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers. Notably, RTD also demonstrates strong cross-classifier transferability: RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify. This confirms that the toxic behaviors induced by RTD are not artifacts of a single classifier but reflect a general vulnerability in VLM alignment under multimodal inputs.

### v4

@@ IV-B Greedy Prompt Search
The greedy prompt search refines these initial prompts iteratively using feedback from the target VLM and a toxicity classifier (e.g., Detoxify [Detoxify]). At each iteration, Gemini proposes a new candidate prompt, from which the diffusion model synthesizes a semantically adversarial image. The toxicity of the resulting VLM continuation is then evaluated, given the generated image and fixed text prefix. If the toxicity score improves over the previous iteration, the new prompt is retained; otherwise, the process terminates. The resulting optimized prompts provide strong initialization points for subsequent reinforcement fine-tuning of the diffusion model.
IV-C Reward Functions for RedDiffuser
After identifying adversarial image prompts through greedy search, we further fine-tune the diffusion model using reinforcement learning guided by a dual reward: a toxicity reward, which encourages harmful continuations, and an alignment reward, which maintains semantic consistency between the image and its image prompt.
Toxicity Reward: Given a fixed text prefix $\mathbf{y}$ and an image $\mathbf{x}$ generated by RedDiffuser, we prompt the target VLM $\mathcal{M}$ to complete the sentence. The continuation is scored by Detoxify [Detoxify], which evaluates six toxicity attributes (e.g., insult, threat). We use the average score as the toxicity reward:
@@ IV-C Reward Functions for RedDiffuser
$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$ (3)
@@
Alignment Reward: To ensure that generated images remain faithful to their image prompts, we introduce an alignment reward inspired by [black2023training]. Specifically, we use BERTScore [zhang2019bertscore] to measure semantic similarity between the image prompt $p$ and the VLM’s interpretation of the generated image:
@@
$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big(S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big)$ (4)
where $\mathcal{M}(\mathbf{x},\text{``Describe the image in short."})$ is the generated image description from the VLM. $S$ calculates the cosine similarity of BERT embeddings between $p_{i}$ and $\mathcal{M}(\mathbf{x},\text{``Describe..."}))_{j}$. $|p|$ is the number of words in the image prompt. This alignment reward acts as a semantic constraint to avoid prompt-irrelevant image drift and model collapse during optimization.
@@
$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$ (5)
This combined reward allows the RedDiffuser to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RedDiffuser is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
Algorithm 1 Training Procedure of RedDiffuser
@@
23:return Fine-tuned RedDiffuser $D_{\theta}^{*}$.
IV-D Extensions for Attacking VLMs with External Guardrails
Real-world deployments often place external guardrails (e.g., image-level NSFW filters) upstream of VLMs, which can block visually explicit images before they influence downstream text generation. To evaluate attacks under this operational constraint, we consider a setting where generated images must both induce toxic continuations and be accepted by external checkers.
Concretely, we adopt a dual-guardrail setup. The first guardrail is the Stable Diffusion image-level NSFW checker; the second is a VLM-based semantic checker implemented by prompting a VLM to judge whether an image contains sexual, violent, or otherwise NSFW content. We treat an image as safe only if both checkers classify it as non-NSFW; images flagged by either checker are refused and do not produce continuations.
To adapt RedDiffuser to this setting, we introduce two complementary extensions: (1) light-weight modifications to the greedy prompt search that incorporate guardrail signals to re-score candidate prompts, and (2) a checker-aware reward shaping mechanism that masks the toxicity reward when a generated image is flagged by the external safety checker.
Greedy Prompt Search with Guardrail Signals
Given the image prompts obtained from the base greedy search, we generate ten images per prompt due to the stochasticity of diffusion sampling. If any sampled image passes both guardrails, the prompt is retained. If all attempts are blocked, we regenerate the image prompt by prompting the attacker LLM with explicit awareness of the external guardrails; the regeneration then follows the same greedy search procedure as before, except that any image prompt that produces NSFW images is skipped. This selective filter-and-regenerate step minimally adapts the prompt set to the guardrail-augmented setting.
Checker-Aware Reward Shaping
When external guardrails are applied, images that fail either check are directly refused. To account for this, we introduce a binary indicator function $\mathbb{I}(\mathbf{x})$ that determines whether a generated image $\mathbf{x}$ passes the dual guardrails:
$\mathbb{I}(\mathbf{x})=\mathbf{1}\{g_{1}(\mathbf{x})=\text{safe}\ \land\ g_{2}(\mathbf{x})=\text{safe}\}$ (6)
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, and $g_{1}$ and $g_{2}$ denote the Stable Diffusion NSFW checker and the VLM-based semantic checker, respectively.
We condition the toxicity reward on this indicator:
$\mathcal{R}_{\text{tox}}^{\text{guard}}(\mathbf{x},\mathbf{y})=\mathbb{I}(\mathbf{x})\cdot\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y})).$ (7)
The alignment reward remains unchanged, giving the total reward:
$\mathcal{R}_{\text{total}}^{\text{guard}}=\mathcal{R}_{\text{tox}}^{\text{guard}}+\lambda\mathcal{R}_{\text{align}}$ (8)
This checker-aware shaping drives RedDiffuser to discover adversarial image styles that are subtle enough to evade guardrails while remaining effective at eliciting toxic continuations.
V Experiments
@@
Datasets We use the RealToxicityPrompts benchmark [gehman2020realtoxicityprompts], specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations [schick2021self, mehrabi2022robust, qi2024visual]. We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RedDiffuser’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RedDiffuser’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
@@
Metrics: (1) Toxicity Rate. To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark [gehman2020realtoxicityprompts], we use the Detoxify classifier [Detoxify] during training, and both Detoxify and the Perspective API during testing. These classifiers assess toxicity across six predefined attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). For evaluation, we compute the proportion of continuations whose toxicity score exceeds a threshold of 0.5 for each attribute. The use of the Perspective API at test time enables cross-classifier validation, helping to verify that the RedDiffuser’s behavior is not overfitted to Detoxify and thus not a result of reward hacking or classifier-specific bias. (2) CLIPScore. To evaluate the semantic alignment between the generated images and the image prompts, we use CLIPScore [hessel2021clipscore], a reference-free metric based on the cosine similarity between CLIP-encoded image and text features. Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following [hessel2021clipscore]) for alignment with human judgment. In our setting, CLIPScore serves as an auxiliary metric to verify that adversarial images preserve semantic plausibility without compromising visual coherence. (3) Guardrail Pass Rate (GPR). When external image-level guardrails are applied, images flagged as NSFW are blocked before producing a continuation. To measure how often generated images remain admissible, we report the Guardrail Pass Rate (GPR), defined as the average value of the indicator in Eq. (6):
$\text{GPR}=\frac{1}{N}\sum_{\mathbf{x}}\mathbb{I}(\mathbf{x})$ (9)
where $N$ is the number of samples and $\mathbb{I}(\mathbf{x})$ indicates whether the image passes both guardrails.
@@
Implementation Details We use Stable Diffusion v1.5 [rombach2022high] as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RedDiffuser is conducted using 8 NVIDIA A100 GPUs, with 6 for training and 2 for reward computation. Reinforcement fine-tuning uses a batch size of 24 per GPU and a learning rate of 3e-4. We apply early stopping after 600 updates to mitigate reward hacking [gao2023scaling]. We experiment with four alignment reward weights $\lambda$: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
@@
TABLE I: Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. Any* indicates the text exhibits at least one of the six toxic attributes.
@@
RedDiff w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
RedDiff w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
RedDiff w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
RedDiff w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
@@
RedDiff w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
RedDiff w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
RedDiff w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
RedDiff w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
V-A Main Results
To evaluate the effectiveness of RedDiffuser (RedDiff), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM. We compare RedDiffuser against several baseline strategies to assess its advantage in inducing harmful outputs. We first include a Text-Only Attack, in which the model is prompted solely with toxic text prefixes without any accompanying image. This serves as a unimodal baseline to quantify the additional effect of visual inputs on toxicity. Next, we define the Single-Prompt Attack, where Gemini generates a single image prompt per prefix without any optimization. This provides a naive multimodal baseline, capturing the effect of directly pairing a non-optimized image with a toxic prefix. The Greedy-Prompt Attack improves upon this by iteratively refining the image prompt using our proposed greedy search algorithm. Finally, we evaluate the full RedDiffuser framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards. We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RedDiff w/ alignment (L)), 0.1 (RedDiff w/ alignment (M)), and 0.2 (RedDiff w/ alignment (H)), as discussed in the implementation details.
TABLE II: Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
@@ V-A Main Results
RedDiff w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
RedDiff w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
RedDiff w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
RedDiff w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
@@
RedDiff w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
RedDiff w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
RedDiff w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
RedDiff w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
@@
Original Attack Set Table I presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the generated outputs, with ‘Any*’ indicating whether a continuation exhibits at least one of the six toxic attributes. We begin by examining the Text-Only Attack, where only the toxic text prefix is provided without any visual input. Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate, achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective API. This indicates that current alignment mechanisms in VLMs remain insufficient for preventing harmful completions even in unimodal settings, exposing a fundamental vulnerability in their default safety alignment. When introducing visual input, the Single-Prompt Attack slightly alters the toxicity pattern but does not significantly improve attack effectiveness, serving as a weak multimodal baseline. In contrast, the Greedy-Prompt Attack significantly amplifies toxicity. For instance, using the Detoxify classifier, the ‘Any*’ rate increases from 55.72% (Single-Prompt) to 64.12%, illustrating the benefit of greedy prompt search in finding high-risk image prompts. The full RedDiffuser framework further enhances attack strength. RedDiff w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers. Notably, RedDiffuser also demonstrates strong cross-classifier transferability: RedDiff w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify. The consistent gains across Detoxify and Perspective API indicate that RedDiffuser’s induced toxic behaviors are not tied to a specific classifier, but instead exploit broader alignment vulnerabilities in VLMs.
Hold-out Set Table II presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed RedDiffuser. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy-Prompt Attack demonstrates a clear improvement over the Single-Prompt Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Single-Prompt Attack. Similarly, with the Perspective API, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Single-Prompt Attack. More notably, the full RedDiffuser framework consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RedDiff w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RedDiff w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These results confirm that RedDiffuser generalizes well to unseen prompts and performs consistently across toxicity classifiers, highlighting its effectiveness as a black-box multimodal red teaming framework.
V-B Cross-Model Transferability
We assess the transferability of our method by applying RedDiffuser trained solely using feedback from LLaVA [liu2024visual] to two unseen VLMs: Gemini-1.5-flash [team2023gemini] and LLaMA-3.2-11B-Vision-Instruct. Tables III and IV report the toxicity rates measured by both Detoxify and the Perspective API. RedDiffuser maintains high toxicity across models despite never interacting with them during training, demonstrating strong generalization.